OpenStudy (anonymous):

seven less than 4 times the square of a number is 18. Find the number.

OpenStudy (anonymous):

Well first write the equation. 4x^2-7=18 Add 7 to both sides. 4x^2-7+7 = 18 +7 4x^2 = 25 Divide all sides by 4. 4x^2/4 = 25/4 x^2=6.25 x=2.5